Year Cash Flow -$7,700 5,600 5,800 4,200 Required: (a) What is the profitability index for the cashflows if the relevant discount rate is 10 percent? (Click to select (b) What is the profitability index for the cashflows if the relevant discount rate is 16 percent? (Click to select) (c) What is the profitability index for the cashflows if the relevant discount rate is 23 percent? (Click to select)
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

At 10%:

Present value of inflows=5600/1.1+5800/1.1^2+4200/1.1^3

=$13039.82

PI=Present value of inflows/Present value of outflows

=$13039.82/$7700

=1.69(Approx).

At 16%:

Present value of inflows=5600/1.16+5800/1.16^2+4200/1.16^3

=$11828.69

PI=Present value of inflows/Present value of outflows

=$11828.69/7700

=1.54(Approx)

At 23%:

Present value of inflows=5600/1.23+5800/1.23^2+4200/1.23^3

=$10643.55

PI=Present value of inflows/Present value of outflows

=$10643.55$7700

=1.38(Approx).